OpenStudy (esraa):

Two parallel plate capacitors, C1 and C2 are connected in series to a 9.0-V battery. If capacitor C1 has a capacitance of 10 nF and capacitor C2 has a capacitance of 40 nF, what is the magnitude of the charge on each plate of capacitor C2? Two parallel plate capacitors, C1 and C2 are connected in series to a 9.0-V battery. If capacitor C1 has a capacitance of 10 nF and capacitor C2 has a capacitance of 40 nF, what is the magnitude of the charge on each plate of capacitor C2?

OpenStudy (marco26):

Hmmm..I think this is how it will go: Since the arrangement is in series, \[qtotal = q1 =q2\] and\[V= V1 + V2\] First, get the Ceq or the C equivalent:\[\frac{1}{\frac{1}{10}+\frac{1}{40}}= 8nf\] \[Ceq = \frac{qtot}{V}\]\[qtot = CeqV = 8(9) = 72 nC\] That is is the magnitude of the charge on each plate of capacitor C2, as well as of C1

OpenStudy (esraa):

Thnx alot =)